# A Survey on Sequential Recommendation

### v1

@@ Abstract
Different from most conventional recommendation problems, sequential recommendation focuses on learning users’ preferences by exploiting the internal order and dependency among the interacted items, which has received significant attention from both researchers and practitioners. In recent years, we have witnessed great progress and achievements in this field, necessitating a new survey. In this survey, we study the SR problem from a new perspective (i.e., the construction of an item’s properties), and summarize the most recent techniques used in sequential recommendation such as pure ID-based SR, SR with side information, multi-modal SR, generative SR, LLM-powered SR, ultra-long SR and data-augmented SR. Moreover, we introduce some frontier research topics in sequential recommendation, e.g., open-domain SR, data-centric SR, could-edge collaborative SR, continuous SR, SR for good, and explainable SR. We believe that our survey could be served as a valuable roadmap for readers in this field.
@@ 1 Introduction
Various sequential recommendation (SR) models have been proposed, which have achieved significant improvements in performance recently [1]. The main idea of an SR model is to exploit the position or sequential information of the interacted items, so as to capture the dependencies among them. Given a user’s interaction sequence, a typical SR model aims to predict the next likely-to-be-preferred items for this user. In recent years, there has been a significant increase in the number of SR models. By searching the keywords (i.e., sequential, recommend) in the DBLP website111https://dblp.org/, we have the specific publication numbers of papers about sequential recommendation, which are shown in Figure 1. We can see that the number of papers about sequential recommendation increases with time going by. Note that the number of papers published in 2024 is slightly fewer than in 2023, because we collected the data on June 12th, 2024.
@@
Some early SR models [2, 1] use a unique ID to denote an item. These SR models are simple yet effective. For each item, an SR model usually needs to learn a unique item embedding, which is often efficient. Meanwhile, the item embeddings can capture the correlations among the items and extract their latent characteristics. However, these SR models have some weaknesses. For instance, only using item IDs cannot solve cold-start and sparsity problems. Therefore, some SR models [3, 4] combine item IDs and item features. The item features mainly include categorical features (e.g., categories), numerical features (e.g., prices) and graph structure features (e.g., social networks and knowledge graph). Due to privacy issues, we often know little about the users. By leveraging these features, we can effectively learn item representations even when interactions with some items are sparse. Therefore, these SR models can alleviate the cold-start and data sparsity problems. However, these SR models also have some limitations. For example, they usually heavily rely on item IDs during training. If data are different, the item IDs are different. Therefore, the models trained on one dataset cannot be easily transferred to another dataset through fine-tuning or other methods. With the fast development of natural language processing and computer vision, models like BERT [5] and ResNet [6] can extract rich features from text descriptions and images. To solve the issue of data dependency, some SR models [7, 8] leverage BERT and ResNet to extract semantic representations of items from text descriptions and images, respectively, and then directly use them to represent the items. Due to the widespread availability of multi-modal features, these models trained on one data from a certain platform can be deployed on another platform through fine-tuning. However, the semantic representations extracted from multi-modal features cannot fully capture users’ fine-grained preferences and collaborative signals. Therefore, some SR models turn to combine the multi-modal features with item IDs to learn users’ preferences more comprehensively.
@@
LLMs are trained using a large amount of data and demonstrate promising performance. More importantly, LLMs have a strong reasoning ability. Therefore, LLMs have been applied to many applications, including recommender systems. For example, some SR models leverage LLMs to recommend items to users directly [9]. Some SR models leverage the output representations of LLMs for recommending personalized items more accurately since the embeddings contain rich semantic knowledge [10]. To alleviate data sparsity and cold-start problems, some SR models leverage LLMs to generate some data. These SR models then combine the generated data with the original data to achieve better recommendation performance. However, LLMs also have some shortcomings. For example, training LLMs-based SR models requires significant computational resources. This survey summarizes how to leverage LLMs in SR models.
@@
In sequential recommendation, modeling ultra-long interaction sequences becomes increasingly complex. Over time, users interact with more and more items. With the length of interaction sequences increasing, the time of both training and inference will increase. Meanwhile, as users interact with an increasing number of items, noise also becomes more serious in the interaction sequences. Therefore, modeling ultra-long interaction sequences will become increasingly difficult. Common approaches retrieve a few interacted items from ultra-long interaction sequences [11]. Our surveys will summarize some common methods applied to ultra-long interaction sequences modeling.
@@
In our survey, we also introduce data augmentation methods in SR models. This is because data augmentation methods can be leveraged by different SR models. They mainly focus on generating more interaction data to solve data sparsity and cold-start problems. Common augmentation methods include some operations, such as resorting, masking, cropping, and others on interaction sequences [12].
@@
Our survey summarizes the latest techniques applied to sequential recommendation
@@ 2 Related Surveys
In recent years, an increasing number of surveys on sequential and session-based recommender systems have been published [13, 14, 15, 16, 17, 18]. They introduce some common algorithms and models for these systems. However, with the development of LLMs and other techniques, these surveys are out-of-date to some extent as LLMs have strong deduction abilities and can fully extract semantic information for recommendation. Meanwhile, some SR models introduced in these surveys still have weaknesses. For example, only using an item ID to denote an item might influence the transferability of SR models. Some surveys [19, 20] focus on cross-domain sequential recommender systems, or multi-behavior sequential recommender systems. They are only a small fraction of models in SR models. Compared with these surveys, our survey introduces the development of sequential recommender systems more comprehensively. There are also some surveys about recommender systems. Jing et al. [21] introduce contrastive learning in recommender systems comprehensively. Lai et al. [22] discuss the importance of data and some methods of data augmentation. Zhang et al. [23] explore the usage of different kinds of side information in recommender systems. However, these surveys only focus on one specific certain technique in recommender systems. Compared with these surveys, we not only classify existing research on SR but also introduce a new taxonomy to comprehend, structure and analyze these works.
@@
The item denoted by tokens | |
@@ 3.2 Properties
In sequential recommendation systems, there are two important entities (i.e., an item and a user). Therefore, item properties and user properties are vital for SR models. A user interacts with items, and the user and the items are connected according to behaviors. The user’s interaction sequence is constructed according to interaction order. Therefore, the properties of both behaviors and sequences are also important in sequential recommendation systems. The specific details about these properties in sequential recommendation systems are shown in Figure 2.
@@
Entity IDs (i.e., item IDs and user IDs), general features and multi-modal features are quietly different in sequential recommender systems. The entity IDs are specific to a particular dataset. In different datasets or different platforms, the encoding methods for the entity IDs are quite different. For example, a movie might be encoded as item in the MovieLens dataset 222https://grouplens.org/datasets/movielens/. However, a book might also be encoded as item 1 in the book domain of Amazon dataset 333https://huggingface.co/datasets/McAuley-Lab/Amazon-Reviews-2023. Therefore, item embeddings based on item IDs cannot be transferred from one platform to another platform. The general features share some common characteristics across different datasets or platforms. For example, basic behavior types in different datasets might be the same. All e-commerce platforms contain basic behavior types (i.e., click and purchase). However, the categories of items might differ across different domains or platforms, leading to different distributions of general features. If we apply the embeddings of general features in one domain to another domain, it might result in negative transfer. Techniques like a cross-attention mechanism can address this issue to some extent. Multi-modal features, such as text, images, and videos, are universal across all domains and platforms. We use language to record information as text, cameras to capture beautiful images, videos to document moments. The text, images, and videos have separate embedding spaces. Meanwhile, they all have shared space. The multi-features can be leveraged in all domains and all platforms. If we leverage the text descriptions of an item to represent the item in a model and train the model on one platform, it can be directly applied to another platform or fine-tuned as needed. Therefore, leveraging multi-modal features to denote an item can fully realize the knowledge transfer from one platform to another.
In the meantime, the general features and multi-modal features can be converted into each other. As shown in Figure 3, we can convert an item’s brand, price and color into text descriptions. Then, SR models can leverage these text descriptions of this item to learn the universal representations of this item. However, this type of conversion has shortcomings. For example, if we convert timestamps to text, fine-grained time information might be damaged. Meanwhile, as shown in Figure 3, we can also extract general features from the text descriptions of an item, such as its brand, color and price. However, the extracted features might contain some noise.
@@
Image refers to the visual representations of an item. It can provide a clear and intuitive understanding of the item. By looking at images, we can know what the item looks like.
Rating refers to the degree of a user’s satisfaction with an item. It can reflect the user’s preferences for an item and the quality of the item.
Review refers to comments that a user provides for an item. It can offer more detailed information about a user’s preferences for the item, and it can reflect the overall quality of this item.
@@ 3.3 Categories
In our surveys, we classify SR models into four categories based on the properties of items. As depicted in Figure 4, for user , there are three different interaction sequences containing item IDs and different features, respectively.
@@
As shown in part (a) of Figure 4, some early works leverage a unique item ID to denote an item for recommendation [1, 24]. These SR models leverage only item ID interaction sequences, which can be formalized as: . In these works, given a candidate item and a user ’s historical interaction sequence , these models need to calculate the probability in which user interacts with the candidate item .
@@
However, SR models based on pure item IDs cannot tackle cold-start problems well. General features are important for recommending personalized items, which can alleviate the cold-start problems and fully extract users’ preferences. As depicted in part (a) and part (c) of Figure 4, some works leverage both item ID sequences and interaction sequences containing general features (i.e., categorical features and numerical features) to recommend personalized items to users [25, 4]. These SR models can be formalized as: . Given a user ’s interaction sequence , which contains item IDs and general features embedded in the sequence, these models aim to predict the probability that the user interacted with a candidate item . To account for the influence of users’ friends, social networks might be leveraged in these works [26]. The general features of interacted items can be represented as knowledge graph. Therefore, these works might leverage knowledge graph to capture relationships among general features.
@@
Though the above-mentioned models achieve significant performance, they still have certain shortcomings. Firstly, these models heavily rely on item IDs, which hinders their transferability. Secondly, these general features cannot utilize some of latest techniques effectively, such as LLMs. Therefore, some recently proposed models leverage the multi-modal features of an item to denote the item directly [8, 27]. These models can realize the transferability across different platforms. By leveraging latest techniques like LLMs, they can fully extract some semantic knowledge. These models can be formalized as: . Given the multi-modal features derived from the items within a user ’s interaction sequence, these models can calculate the probability in which user interacts with a candidate item . The core ideas of these models are illustrated in part (b) of Figure 4. When both sequence information and multi-modal features (i.e., text descriptions and images) of each interaction are available, an SR model predicts the next interacted item interacted with by user . By replacing item IDs with text descriptions or images, SR models can require rich semantic knowledge and enhance their transferability.
@@
However, relying solely on items’ multi-modal features cannot capture users’ fine-grained interests. In a warm-start situation, only leveraging item IDs can achieve better performance than only leveraging multi-modal features [7]. Therefore, some SR models leverage both item IDs and multi-modal features of these items. The simple diagram of these models is shown by combining part (a) and part (b) in Figure 4. In these SR models, each item property contains both a unique item ID and multi-modal features. To extract semantic information from multi-modal features, LLMs can be leveraged. These SR models can be formalized as: .
@@ 4 Pure ID-based SR
In SR models based on pure IDs, each item and user are represented by a unique item identity (i.e., item ID) and a unique user identity (i.e., user ID), respectively. In the past few years, the number of SR models based on pure IDs has increased significantly. We categorize them into different classes according to the used techniques. In this section, we will first introduce some SR methods using traditional machine learning. Then, we will introduce some SR methods integrating various deep learning techniques. Next, we will introduce some SR methods using reinforcement learning. Finally, we will introduce some SR methods that combine different kinds of deep learning techniques. In this section, all SR models are based on pure IDs.
@@ 4.1 Traditional Models
FPM (frequent pattern mining) extracts different kinds of patterns from interaction sequences [28]. These models are of good explainability, but they cannot tackle complex datasets. If the dataset is too large, the number of extracted sequential patterns might become excessively too large.
Markov models have an assumption that a predicted action relies on a few recent actions. For example, FPMC [29] predicts the next interacted item for each user by considering both the last interacted item and personalized interests. These models cannot exploit long-term sequential dependencies. If order dependencies are weak in a dataset, Markov models might perform poorly.
Latent factor models mainly learn latent representations to estimate unobserved transitions. TransRec [2] proposes a transition operator: a previously interacted item + a user the user next interacted item. The transition operator might vary across different datasets. Meanwhile, these SR models cannot capture complex transition relations.

### v2

@@ Abstract
Different from most conventional recommendation problems, sequential recommendation focuses on learning users’ preferences by exploiting the internal order and dependency among the interacted items, which has received significant attention from both researchers and practitioners. In recent years, we have witnessed great progress and achievements in this field, necessitating a new survey. In this survey, we study the SR problem from a new perspective (i.e., the construction of an item’s properties), and summarize the most recent techniques used in sequential recommendation such as multi-modal SR, generative SR, LLM-powered SR, ultra-long SR and data-augmented SR. Moreover, we introduce some frontier research topics in sequential recommendation, e.g., open-domain SR, data-centric SR, cloud-edge collaborative SR, continuous SR, SR for good, and explainable SR. We believe that our survey could be served as a valuable roadmap for readers in this field.
@@ 1 Introduction
Various sequential recommendation (SR) models have been proposed, which have achieved significant recommendation improvements in performance recently [1]. The main idea of an SR model is to exploit the position and sequential information of the interacted items, so as to capture the dependencies among them. Given a user’s interaction sequence, a typical SR model aims to predict the next likely-to-be-preferred items for this user. In recent years, there has been a significant increase in the number of SR models. By searching the keywords (i.e., sequential, recommend) in the DBLP website111https://dblp.org/, we have the specific publication numbers of papers about sequential recommendation, which are shown in Figure 1. We can see that the number of papers about sequential recommendation increases with time going by. Note that the number of papers published in 2024 is slightly fewer than in 2023, because we collected the data on June 12th, 2024.
@@
Some early SR models [2, 1] use a unique ID to denote an item. These SR models are simple yet effective. For each item, an SR model usually needs to learn a unique item embedding, which is often efficient. Meanwhile, the item embeddings can capture the correlations among the items and extract their latent characteristics. However, these SR models have some weaknesses. For instance, only using item IDs cannot solve cold-start and sparsity problems. Therefore, some SR models [3, 4] combine item IDs and item features. The item features mainly include categorical features (e.g., categories), numerical features (e.g., prices) and graph structure features (e.g., social networks and knowledge graph). Due to privacy issues, we often know little about the users’ features. By leveraging these features, we can effectively learn item representations even when interactions with some items are sparse. Therefore, these SR models can alleviate the cold-start and data sparsity problems. However, these SR models also have some limitations. For example, they usually heavily rely on item IDs during training. If data are different, the item IDs are different. Therefore, these SR models trained on one data cannot be easily transferred to another data through fine-tuning or other methods. With the fast development of natural language processing and computer vision, models like BERT [5] and ResNet [6] can extract rich features from text descriptions and images, respectively. To solve the issue of data dependency and high-difficulty in cross-data transfer, some SR models [7, 8] leverage BERT and ResNet to extract semantic representations of items from text descriptions and images, respectively, and then directly use them to represent the items. Due to the widespread availability of multi-modal features, these SR models trained on one data from a certain platform can be deployed on another platform through fine-tuning. However, the semantic representations extracted from multi-modal features cannot fully capture users’ fine-grained preferences and collaborative signals. Therefore, some SR models turn to combine the multi-modal features with item IDs to learn users’ preferences more comprehensively.
@@
LLMs are trained using a large amount of data and demonstrate promising performance. More importantly, LLMs have a strong reasoning ability. Therefore, LLMs have been applied to many applications, including recommender systems. For example, some SR models leverage LLMs to recommend items to users directly [9]. Some SR models leverage the output representations of LLMs for recommending personalized items more accurately since these representations contain rich semantic knowledge [10]. To alleviate data sparsity and cold-start problems, some SR models leverage LLMs to generate some data. These SR models then combine the generated data with the original data to achieve better recommendation performance. However, LLMs-based SR models also have some shortcomings. For example, training LLMs-based SR models requires significant computational resources. This survey summarizes how to leverage LLMs in SR models.
@@
In sequential recommendation, modeling ultra-long interaction sequences becomes increasingly complex. Over time, users interact with more and more items. With the length of interaction sequences increasing, the time of both training and inference will increase. Meanwhile, as users interact with an increasing number of items, noise also becomes more serious in the interaction sequences. Therefore, modeling ultra-long interaction sequences becomes increasingly difficult. Common approaches retrieve a few interacted items from ultra-long interaction sequences [11]. Our survey will summarize some common methods applied to modeling ultra-long interaction sequences modeling.
@@
In our survey, we also introduce data augmentation methods in SR models. This is because data augmentation methods can be leveraged by different SR models. They mainly focus on generating more interaction data to alleviate data sparsity and cold-start problems. Common augmentation methods include some operations, such as resorting, masking, cropping, and others on interaction sequences [12].
@@
Our survey summarizes the latest techniques applied to sequential recommendation.
@@
The remainder of our survey is organized as follows. In Section 2, we discuss some surveys on sequential recommendation and session-based recommendation. In Section 3, we provide a comprehensive view about sequential recommendation, including problem definition, properties, four different kinds of SR models and challenges. In Section 4, we discuss some pure ID-based SR models. In Section 5, we present some SR models combining IDs and side information. In Section 6, we elaborate on recent advancements in sequential recommendation, including multi-modal SR, generative SR, LLM-powered SR, ultra-long sequence modeling in SR, and data augmentation in SR. In Section 7, we introduce datasets, evaluation protocols and metrics, as well as experimental results in sequential recommendation. Finally, we provide some insights into prospects and future directions about sequential recommendation in Section 8 and draw a conclusion in Section 9.
@@ 2 Related Surveys
In recent years, an increasing number of surveys on sequential and session-based recommender systems have been published [13, 14, 15, 16, 17, 18]. They introduce some common algorithms and models for these systems. However, with the development of LLMs and other techniques, these surveys are out-of-date to some extent as LLMs have strong deduction abilities and can fully extract semantic information for recommendation. Meanwhile, some SR models introduced in these surveys still have weaknesses. For example, only using an item ID to denote an item might influence the transferability of SR models. Some surveys [19, 20] focus on cross-domain sequential recommender systems, or multi-behavior sequential recommender systems. They only introduce a small fraction of models in sequential recommendation. Compared with these surveys, our survey introduces the development of sequential recommender systems more comprehensively. There are also some surveys about recommender systems. Jing et al. [21] introduce contrastive learning in recommender systems comprehensively. Lai et al. [22] discuss the importance of data and some methods of data augmentation. Zhang et al. [23] explore the usage of different kinds of side information in recommender systems. However, these surveys only focus on one specific technique in recommender systems. Compared with these surveys, we not only classify existing research on SR but also introduce a new taxonomy to comprehend, structure and analyze these works.
@@
| The tokens representing item | |
@@ 3.2 Properties
In sequential recommendation systems, there are two important entities (i.e., an item and a user). Therefore, item properties and user properties are vital for SR models. A user interacts with items, establishing connections between the user and the items via different types of behaviors. The user’s interaction sequence is constructed according to interaction order. Therefore, the properties of both the behaviors and sequences are also important in sequential recommendation systems. The specific details about these properties in sequential recommendation systems are shown in Figure 2.
@@
Entity IDs (i.e., item IDs and user IDs), general features and multi-modal features are quietly different in sequential recommender systems. The entity IDs are specific to a particular data. In different data or different platforms, the encoding methods for the entity IDs are quite different. For example, a movie might be encoded as item in MovieLens 222https://grouplens.org/datasets/movielens/. However, a book might also be encoded as item 1 in the book domain of Amazon 333https://huggingface.co/datasets/McAuley-Lab/Amazon-Reviews-2023. Therefore, item embeddings based on item IDs cannot be transferred from one platform to another platform. The general features share some common characteristics across different datasets or platforms. For example, basic behavior types in different datasets might be the same. All e-commerce platforms contain basic behavior types (i.e., click and purchase). However, the categories of items might differ across different domains or platforms, leading to different distributions of general features. If we apply the embeddings of general features in one domain to another domain, it might result in negative transfer. Techniques like a cross-attention mechanism can address this issue to some extent. Multi-modal features, such as text, images, and videos, are universal across all domains and platforms. We use language to record information as text, cameras to capture beautiful images, videos to document moments. The text, images, and videos have separate embedding spaces. Meanwhile, they all have shared space. Therefore, the multi-features can be leveraged in all domains and all platforms. If we leverage the text descriptions of an item to represent the item in a model and train the model on one platform, it can be directly applied to another platform or fine-tuned in another platform as needed. Therefore, leveraging multi-modal features to denote an item can fully realize the transfer of SR models from one platform to another.
In the meantime, the general features and multi-modal features can be converted into each other. As shown in Figure 3, we can convert an item’s brand, price and color into text descriptions. Then, SR models can leverage these text descriptions of this item to learn the universal representations of this item. However, this type of conversion has shortcomings. For example, if we convert timestamps to text, fine-grained time information might be damaged. Meanwhile, as shown in Figure 3, we can also extract general features from the text descriptions of an item, such as its brand, color and price. However, the extracted general features might contain some noise.
@@
Image refers to the visual representations of an item. It can provide a clear and intuitive understanding of the item. By looking at images, users can know what the item looks like.
Rating refers to the degree of a user’s satisfaction with an item. It can reflect the user’s preferences for the item and the quality of the item.
Review refers to comments that a user provides for an item. It can offer more detailed information about the user’s preferences for the item, and it can reflect the overall quality of this item.
@@ 3.3 Categories
In our surveys, we classify SR models into four categories based on the properties of items. As depicted in Figure 4, for user , there are three different interaction sequences containing item IDs, multi-modal features, and categorical and numerical features, respectively.
@@
As shown in part (a) of Figure 4, some early works leverage a unique item ID to denote an item for recommendation [1, 24]. These SR models leverage only item ID interaction sequences, which can be formalized as: . In these works, given a candidate item and a user ’s historical interaction sequence , these models need to calculate the probability in which user interacts with the candidate item . We will introduce this type of SR models in Section 4.
@@
However, SR models based on pure item IDs cannot tackle the commonly encountered cold-start and data-sparsity problems well. General features are important for recommending personalized items, which can alleviate the cold-start and data-sparsity problems and fully extract users’ preferences. As depicted in part (a) and part (c) of Figure 4, some works leverage both item ID sequences and interaction sequences containing general features (i.e., categorical features and numerical features) to recommend personalized items to users [25, 4]. These SR models can be formalized as: . Given a user ’s interaction sequence , which contains item IDs and general features embedded in the sequence, these models aim to predict the probability in which the user interacted with a candidate item . To account for the influence of users’ friends, social networks might be leveraged in these works [26]. The general features of interacted items can be represented as knowledge graph. Therefore, these works might leverage knowledge graph to capture relationships among general features. We will introduce this type of SR models in Section 5.
@@
Though the above-mentioned models achieve significant performance, they still have certain shortcomings. Firstly, these models heavily rely on item IDs, which hinders their transferability. Secondly, these general features might not be fully utilized by some of latest techniques effectively, such as LLMs. Therefore, some recently proposed models leverage the multi-modal features of an item to denote the item directly [8, 27]. These models can realize the transferability across different platforms. By leveraging latest techniques like LLMs, they can fully extract some semantic knowledge. These models can be formalized as: . Given the multi-modal features derived from the items within a user ’s interaction sequence, these models can calculate the probability in which user interacts with a candidate item . The core ideas of these models are illustrated in part (b) of Figure 4. When both sequence information and multi-modal features (i.e., text descriptions and images) of each interaction are available, an SR model predicts the next interacted item interacted with by user . By replacing item IDs with text descriptions and images, SR models can acquire rich semantic knowledge and enhance their transferability.
@@
However, relying solely on items’ multi-modal features cannot capture users’ fine-grained interests. In a warm-start situation, only leveraging item IDs can achieve better performance than only leveraging multi-modal features [7]. Therefore, some SR models leverage both item IDs and multi-modal features of these items. The simple diagram of these models is shown by combining part (a) and part (b) in Figure 4. In these SR models, each item property contains both a unique item ID and multi-modal features. To extract semantic information from multi-modal features, LLMs can be leveraged. These SR models can be formalized as: . We will introduce the aforementioned two types of SR models in Section 6.1 of Multi-Modal SR.
@@ 4 Pure ID-based SR
In SR models based on pure IDs, each item and user are represented by a unique item identity (i.e., item ID) and a unique user identity (i.e., user ID), respectively. In the past few years, the number of SR models based on pure IDs has increased significantly. We categorize them into different classes according to the used techniques. In this section, we will first introduce some SR methods using traditional machine learning. Then, we will introduce some SR methods integrating various deep learning techniques. Next, we will introduce some SR methods using reinforcement learning. Finally, we will introduce some SR methods that combine different kinds of deep learning techniques. In this section, all SR models are designed based on pure IDs.
@@ 4.1 Traditional Models
FPM (frequent pattern mining) extracts different kinds of patterns from interaction sequences [28]. These models are of good explainability, but they cannot tackle complex data. If a data is too large, the number of extracted sequential patterns might become excessively too large.
Markov models have an assumption that a predicted action relies on a few recent actions. For example, FPMC [29] predicts the next interacted item for each user by considering both the last interacted item and personalized interests. However, these models cannot exploit long-term sequential dependencies. If the order dependencies are weak in data, Markov models might perform poorly.
Latent factor models mainly learn latent representations to estimate unobserved transitions. TransRec [2] proposes a transition operator: a previously interacted item + a user the user next interacted item. However, the transition operator might vary across different data. Meanwhile, these SR models cannot capture complex transition relations.